The activation energy for the gas phase isomerization of dimethyl maleate is 111 kJ. cis-CH 00CCH-CHCOOCH trans-CH,OOCCH CHCOOCH The rate constant at 693 K is 5.70 10 /s. The rate constant will be 8.49 103 /s at
Answer #1

Ans is 806 K

The inteynakedonmofthis equaion is ki biven data :8-49810 T1 = 693k 4 | KI = 5.70X104 /s 2.0 23x 10 k2= 8.44 × 10-3/s -2.02